OpenStudy (anonymous):

Tina would like to melt down her existing gold jewelry to make a 20 ounce gold necklace that is 90% pure gold. After melting her jewelry, she is left with two batches. One batch is 75% pure. The other batch is 96% pure. How much of the 96% pure melted gold should Tina mix with the 75% pure melted gold to obtain 20 ounces of 90% pure gold?

OpenStudy (anonymous):

@Luigi0210 @iGreen @CO_oLBoY

OpenStudy (mathmate):

weight of 96% batch : weight of 75% batch = 90-75 : 96-90 = 15:6 =5:2 Alternatively, let x=fraction of 96%, and 1-x = fraction of 75% Then solve for x in 0.96x+0.75(1-x)=0.9 0.21x=0.15 x=5/7 So 5/7 of 96%, and 2/7 of 75% (same as 5:2)

OpenStudy (anonymous):

thanks!

OpenStudy (mathmate):

You're welcome! :)